Consider a firm with an EBIT of $1,017,000. The firm finances its assets with $4,840,000 debt (costing 7.7 percent) and 217,000 shares of stock selling at $1700 per share. To reduce risk associated with this financial leverage, the firm is considering reducing its debt by $2,720,000 by selling additional shares of stock. The firm is in the 40 percent tax bracket. The change in capital structure will have no effect on the operations of the firm. Thus, EBIT will remain at $1,017,000. Calculate the EPS before and after the change in capital structure and indicate changes in EPS. (Round your answers to 2 decimal places.) EPS before EPS after Difference
Answer #1

Calculating original EPS:

Total debt = $4,840,000
Interest rate = 7.7%

  • EBIT = $1,017,000
  • Interest expense = 7.7% of $4,840,000 = $372,680
  • EBT = $644,320
  • Taxes (40%) = $257,728
  • Net income = $386,592

Total shares outstanding = 217,000

  • EPS = $386,592 / 217,000 = $1.78

Now, there are 160,000 (= $2,720,000 / $17) more shares issued, so:

Total number of shares = 217,000 + 160,000 = 377,000
Debt has been reduced to $4,840,000 - $2,720,000 = $2,120,000

Calculating final EPS:

  • EBIT = $1,017,000
  • Interest expense = 7.7% of $2,120,000 = $163,240
  • EBT = $853,760
  • Taxes (40%) = $341,504
  • Net income = $512,256

Total shares outstanding = 377,000

  • EPS = $512,256 / 377,000 = $1.36

Difference = EPS before - EPS after = $1.78 - $1.36 = $0.42